OpenStudy (anonymous):

write the expression of either sine cosine or tangent of a single angle cos(pi/5)cos(pi/7)+sin(pi/5)sin(pi/7)

myininaya (myininaya):

do you recall this identity: \[\cos(a-b)=\cos(a)\cos(b)+\sin(a)\sin(b)?\]

myininaya (myininaya):

do you see how you can use that identity here?

OpenStudy (anonymous):

kind of yes

myininaya (myininaya):

is it obvious to you what a and b is? if so just replace a and b on the left hand side

OpenStudy (anonymous):

5 and 7 correct?

myininaya (myininaya):

i don't see 5 and 7

myininaya (myininaya):

i see the following numbers pi/5 and pi/7

OpenStudy (anonymous):

yes thats what i meant!

myininaya (myininaya):

so it doesn't really matter what you put as a and what you put as b since cos is an even function cos(a-b)=cos(-(a-b)) by the fact cosine is even =cos(b-a)

OpenStudy (anonymous):

when I plugged in a and b i got 1

myininaya (myininaya):

so pi/5-pi/7=1? I don't think so unless I have forgotten how to add/subtract fractions

myininaya (myininaya):

you need to find a common denominator

myininaya (myininaya):

what is the lcm (least common multiple) of 5 and 7?

OpenStudy (anonymous):

35

myininaya (myininaya):

ok we have \[\frac{\pi}{5}-\frac{\pi}{7}\] and we need the same denominator which we determined that will be 35 so we have \[\frac{7 \pi}{5 \cdot 7}-\frac{5 \pi}{7 \cdot 5}\] first fraction we multiplied by 7/7 second fraction we multiplied by 5/5

myininaya (myininaya):

now what is 7pi-5pi?

OpenStudy (anonymous):

2pi

myininaya (myininaya):

so pi/5-pi/7 is 2pi/35

myininaya (myininaya):

so \[\cos(\frac{\pi}{5})\cos(\frac{\pi}{7})+\sin(\frac{\pi}{5})\sin(\frac{\pi}{7})=\cos(\frac{\pi}{5}-\frac{\pi}{7})=\cos(\frac{2\pi}{35})\]

OpenStudy (anonymous):

thank you so so much for all of your help!

myininaya (myininaya):

np